### v1

@@ 4 Experimental Setup
We use BRO (Nauman et al., 2024b), an approach based on SAC (Haarnoja et al., 2018) that uses a regularized residual network to represent the Q-values and has been shown to scale well to high capacities. We note that prior work (Nauman et al., 2024b; Lee et al., 2024) showed that scaling width gives better performance that scaling depth in TD-learning. Thus, to study the impact of model size, we vary only the network width in {256, 512, 1024, 2048, 4096}. We consider batch sizes from 4 to 4096 (varied in powers of 2 and UTD ratios of 1, 2, 4, 8. We keep other hyperparameters fixed across all tasks at values suggested by Nauman et al. (2024b). For our initial study, we leverage the results from (Nauman et al., 2024b) on Deepmind Control suite (Tassa et al., 2018). Following prior work (Hansen et al., 2023; Nauman et al., 2024b), we separate these into 7 medium difficulty tasks (referred to as DMC-medium) and 6 hard difficulty tasks (DMC-hard). For these tasks, we fit averages of the tasks for the two suites respectively, building upon the protocol prescribed in Rybkin et al. (2025), to show generalization of our fits across tasks. To further evaluate scaling on harder tasks, we run 4 tasks from DMC and HumanoidBench (Sferrazza et al., 2024), where we make fits for each task individually to show applicability to single tasks. Further details are in Appendix˜B.
@@ 5 Analyzing the Interplay Between Model Size and Batch Size
Rybkin et al. (2025) argues that the best batch size for a given run is a predictable function of the UTD ratio $\sigma$, and often this relationship follows a power law. In simpler terms, this means that the larger the UTD ratio, the smaller the batch size should be for best performance. The conceptual argument behind this relation is an overfitting argument: larger batch sizes increase the risk of overfitting from repeated training. However, this prior analysis holds model size $N$ constant and does not consider its influence on batch size. We extend prior analysis (Rybkin et al., 2025) by considering how model size modulates the effective batch size under fixed UTD ratio, revealing a distinct form of overfitting unique to TD-learning.
Figure 1: Measuring train and validation TD-errors for different batch sizes on h1-crawl. While the training and validation TD-errors reduce as model size increases, for smaller models a larger batch size results in a higher or plateauing final TD-error. This illustrates the role of batch size in modulating overfitting with TD-learning.
@@ 5.2 A Mental Model for TD-Overfitting
To conceptually understand this phenomenon further, consider a low-capacity Q-network. Due to limited representational capacity, such a model will inevitably entangle the features used to predict Q-values across different state-action pairs (Kumar et al., 2022, 2021a). Reducing TD-error on the training data and updating the target network can now change target values on unseen transitions, potentially increasing validation TD-error with every TD update. Moreover, the more accurate and directed the gradient update on the training error, the worse the validation TD-error can get (Figure 2(b, c); full curves in Figure˜12). Larger batch sizes reduce variance produce directed gradients that exacerbate this problem, as fitting the targets for some transitions comes at the expense of others when the model capacity is low.
@@ 5.3 The Role of Dynamic Programming in TD-Overfitting
In Figure 4, we observe that when the TD-targets are generated by a low-capacity Q-network, the passive critic — regardless of its own capacity — exhibits higher validation TD-error. This supports the main insight from our mental model: the TD overfitting phenomenon is driven primarily by the poor generalization of TD-targets produced by low-capacity networks. While increasing the passive critic’s width improves its ability to fit lower quality targets (e.g., a width = 2048 passive critic fits width = 512 targets slightly better than a smaller passive critic), the validation TD-error can increase or slowly decrease over the course of training (see h1-crawl in Figure˜13). Conversely, when the main Q-network generating the targets is larger (e.g., width = 2048), even smaller passive critics (e.g., width = 512 or 1024) can match resulting TD-targets quite well, and validation error decreases over the course of training (see h1-crawl in Figure˜13). This indicates that a large portion of the overfitting dynamics of TD-learning is governed by the target values, and how they evolve during training, irrespective of the main critic (though the targets themselves depend on the main critic).
@@ 7 Partitioning Compute Optimally Between Model Size and UTD
Implications. Our results show that appropriate choices of UTD and model size improve both sample efficiency and compute utilization. At the same time, we find broad regions of near-equivalent performance: multiple (UTD, model-size) settings perform similarly well, so fully optimizing these hyperparameters is often unnecessary to capture most of the gains (Figure˜19 and Figure˜20). Similarly, while the best configuration is environment-dependent, with some tasks benefiting from larger models to begin learning and others from a higher UTD, scaling the model size paired with a mild increase in UTD is often a good starting point. Our framework makes these trade-offs explicit and provides a principled approach to selecting good values for these hyperparameters.
@@ Appendix B Experiment Details
For our experiments, we use a total of 17 tasks from two benchmarks (DeepMind Control (Tunyasuvunakool et al., 2020) and HumanoidBench (Sferrazza et al., 2024)), listed in Table˜3. As described in Appendix˜C, we normalize our returns to $[0,1000]$; optimal $\pi$ returns are pre-normalized. For HumanoidBench, we report the returns listed by authors as the “success bar,” even though it is possible to achieve a higher return. Our experiments fit $\mathcal{D}_{J}(\sigma,N)$ for 20 normalized performance thresholds $J$, spaced uniformly between $J_{\min}$ and $J_{\max}$, inclusive; 20 is an arbitrary choice that we made so as to obtain useful insights about our method while not overwhelming the reader.
@@ D.1 Batch Size Fits $\tilde{B}(\sigma,N)$
Figure 11: Two-dimensional batch size fit $\tilde{B}(\sigma,N)$ grouped by $\sigma$ and $N$, as a completion to Figure˜5.
@@ D.4 Full TD-error curves
We provide full training and validation TD-error curves in Figure˜12, as a completion to Figure˜1. The summary statistics are marked with ‘X’ and correspond to the points used in Figure˜2.
@@ D.5 Passive Critic Learning Curves
We provide the full validation TD error curves over training in Figure˜13. In these plots the summary statistic is marked with an ‘X’, and we provide Figure˜14 as a completion to Figure˜4.
@@ D.6 Data Efficiency Fits $\mathcal{D}_{J_{\max}}(\sigma,N)$
For the following four tasks, we fit data efficiency using the empirically best data efficiency for performance threshold $J_{\max}$ across batch sizes for each $(\sigma,N)$ setting. In Figures˜15 and 16, we show the fits for multiple values of $J$.
@@
Figure 16: Same as Figure˜15, but instead grouped by model size $N$.
@@ D.7 Optimal Budget Partition
We provide plots analogous to Figures˜7 and 8 in Figures˜17, 18, 19 and 20 for DMC-medium and DMC-hard tasks. These data efficiency fits use the shared exponents $\alpha_{J}$, $\beta_{J}$ method described in Appendix˜B.
As shown in Figures˜8, 19 and 20, however, the optimal UTD and model size for a given budget $\mathcal{F}_{0}$ are unpredictable. We verify that these hyperparameters are fundamentally unpredictable in this setting, running at least 50 seeds per UTD and model size at the fitted batch size, for h1-crawl, in Figure˜21. Despite this, the fit for $\mathcal{C}_{\mathcal{F}_{J}}$ achieves considerably lower uncertainty than in Figure˜7, indicating that there is a large range of “reasonable” hyperparameters corresponding to similar data and compute values.

### v2

@@ 4 Experimental Setup
We use BRO (Nauman et al., 2024b) and SimbaV2 (Lee et al., 2025), approaches based on SAC (Haarnoja et al., 2018) that use a regularized residual network to represent the Q-values and have been shown to scale well to high capacities. These prior works showed that scaling width gives better performance that scaling depth in TD-learning. Thus, to study the impact of model size, we vary only the network width in {256, 512, 1024, 2048, 4096}. We consider batch sizes from 4 to 4096 (varied in powers of 2 and UTD ratios of 1, 2, 4, 8. We keep other hyperparameters fixed across all tasks at values suggested by Nauman et al. (2024b). For our initial study, we leverage the results from (Nauman et al., 2024b) on Deepmind Control suite (Tassa et al., 2018). Following prior work (Hansen et al., 2023; Nauman et al., 2024b), we separate these into 7 medium difficulty tasks (referred to as DMC-medium) and 6 hard difficulty tasks (DMC-hard). For these tasks, we fit averages of the tasks for the two suites respectively, building upon the protocol prescribed in Rybkin et al. (2025), to show generalization of our fits across tasks. To further evaluate scaling on harder tasks, we run 4 tasks from DMC and HumanoidBench (Sferrazza et al., 2024), where we make fits for each task individually to show applicability to single tasks. Further details are in Appendix˜B.
@@ 5 Analyzing the Interplay Between Model Size and Batch Size
Rybkin et al. (2025) argues that the best batch size for a given run is a predictable function of the UTD ratio $\sigma$, and often this relationship follows a power law. In simpler terms, this means that the larger the UTD ratio, the smaller the batch size should be for best performance.
Figure 1: Measuring train and validation TD-errors for different batch sizes on h1-crawl. While the training and validation TD-errors reduce as model size increases, for smaller models a larger batch size results in a higher final TD-error. This illustrates the role of batch size in modulating overfitting with TD-learning.
However, this prior analysis holds model size $N$ constant and does not consider its influence on batch size. We extend prior analysis (Rybkin et al., 2025) by considering how model size modulates the effective batch size under fixed UTD ratio, revealing a distinct form of overfitting unique to TD-learning.
@@ 5.2 A Mental Model for TD-Overfitting
To conceptually understand this phenomenon further, consider a low-capacity Q-network. Due to limited representational capacity, such a model will inevitably entangle the features used to predict Q-values across different state-action pairs (Kumar et al., 2022, 2021a). Reducing TD-error on the training data and updating the target network can now change target values on unseen transitions, potentially increasing validation TD-error with every TD update. Moreover, the more accurate and directed the gradient update on the training error, the worse the validation TD-error can get (Figure 2(b, c); full curves in Figure˜13). Larger batch sizes reduce variance produce directed gradients that exacerbate this problem, as fitting the targets for some transitions comes at the expense of others when the model capacity is low.
@@ 5.3 The Role of Dynamic Programming in TD-Overfitting
In Figure 4, we observe that when the TD-targets are generated by a low-capacity Q-network, the passive critic — regardless of its own capacity — exhibits higher validation TD-error. This supports the main insight from our mental model: the TD overfitting phenomenon is driven primarily by the poor generalization of TD-targets produced by low-capacity networks. While increasing the passive critic’s width improves its ability to fit lower quality targets (e.g., a width = 2048 passive critic fits width = 512 targets slightly better than a smaller passive critic), the validation TD-error can increase or slowly decrease over the course of training (see h1-crawl in Figure˜14). Conversely, when the main Q-network generating the targets is larger (e.g., width = 2048), even smaller passive critics (e.g., width = 512 or 1024) can match resulting TD-targets quite well, and validation error decreases over the course of training (see h1-crawl in Figure˜14). This indicates that a large portion of the overfitting dynamics of TD-learning is governed by the target values, and how they evolve during training, irrespective of the main critic (though the targets themselves depend on the main critic).
@@ 7 Partitioning Compute Optimally Between Model Size and UTD
Implications. Our results show that appropriate choices of UTD and model size improve both sample efficiency and compute utilization. At the same time, we find broad regions of near-equivalent performance: multiple (UTD, model-size) settings perform similarly well, so fully optimizing these hyperparameters is often unnecessary to capture most of the gains (Figure˜20 and Figure˜21). Similarly, while the best configuration is environment-dependent, with some tasks benefiting from larger models to begin learning and others from a higher UTD, scaling the model size paired with a mild increase in UTD is often a good starting point. Our framework makes these trade-offs explicit and provides a principled approach to selecting good values for these hyperparameters.
@@ Appendix B Experiment Details
For our experiments, we use a total of 17 tasks from two benchmarks (DeepMind Control (Tunyasuvunakool et al., 2020) and HumanoidBench (Sferrazza et al., 2024)), listed in Table˜3, with the BRO algorithm and architecture (Nauman et al., 2024b). We additionally use 2 tasks from HumanoidBench (h1-crawl, h1-stand) with SimbaV2 (Lee et al., 2025). As described in Appendix˜C, we normalize our returns to $[0,1000]$; optimal $\pi$ returns are pre-normalized. For HumanoidBench, we report the returns listed by authors as the “success bar,” even though it is possible to achieve a higher return. Our experiments fit $\mathcal{D}_{J}(\sigma,N)$ for 20 normalized performance thresholds $J$, spaced uniformly between $J_{\min}$ and $J_{\max}$, inclusive; 20 is an arbitrary choice that we made so as to obtain useful insights about our method while not overwhelming the reader.
@@ D.1 Batch Size Fits $\tilde{B}(\sigma,N)$
Figure 11: Two-dimensional batch size fit $\tilde{B}(\sigma,N)$ grouped by $\sigma$ and $N$, as a completion to Figure˜5, for the BRO algorithm and architecture (Nauman et al., 2024b)
Grouped by UTD ratio $\sigma$ Refer to caption
Grouped by model size NN Refer to caption
Figure 12: Analogous to Figure˜11, with the SimbaV2 architecture (Lee et al., 2025)
@@ D.4 Full TD-error curves
We provide full training and validation TD-error curves in Figure˜13, as a completion to Figure˜1. The summary statistics are marked with ‘X’ and correspond to the points used in Figure˜2.
@@ D.5 Passive Critic Learning Curves
We provide the full validation TD error curves over training in Figure˜14. In these plots the summary statistic is marked with an ‘X’, and we provide Figure˜15 as a completion to Figure˜4.
@@ D.6 Data Efficiency Fits $\mathcal{D}_{J_{\max}}(\sigma,N)$
For the following four tasks, we fit data efficiency using the empirically best data efficiency for performance threshold $J_{\max}$ across batch sizes for each $(\sigma,N)$ setting. In Figures˜16 and 17, we show the fits for multiple values of $J$.
@@
Figure 17: Same as Figure˜16, but instead grouped by model size $N$.
@@ D.7 Optimal Budget Partition
We provide plots analogous to Figures˜7 and 8 in Figures˜18, 19, 20 and 21 for DMC-medium and DMC-hard tasks. These data efficiency fits use the shared exponents $\alpha_{J}$, $\beta_{J}$ method described in Appendix˜B.
As shown in Figures˜8, 20 and 21, however, the optimal UTD and model size for a given budget $\mathcal{F}_{0}$ are unpredictable. We verify that these hyperparameters are fundamentally unpredictable in this setting, running at least 50 seeds per UTD and model size at the fitted batch size, for h1-crawl, in Figure˜22. Despite this, the fit for $\mathcal{C}_{\mathcal{F}_{J}}$ achieves considerably lower uncertainty than in Figure˜7, indicating that there is a large range of “reasonable” hyperparameters corresponding to similar data and compute values.